# Breaking the Dimensional Barrier: A Pontryagin-Guided Direct Policy Optimization for Continuous-Time Multi-Asset Portfolio

### v1

@@ 3.1 Baseline PG-DPO for the Extended Objective
It is important to understand the role of the intermediate values computed during BPTT in Step 4. As BPTT calculates the policy gradient $\nabla J^{(i)}(\dots)$ via the chain rule, it inherently computes the partial derivatives of the realized reward $J^{(i)}(\dots)$ with respect to the intermediate states along path $i$:
@@
These values, $\lambda_{k}^{(i)}$ and $\bm{\eta}_{k}^{(i)}$, are unbiased estimates of the Pontryagin costates associated with the conditional expected utility $J(t_{0}^{(i)},\dots)$. Crucially, in the baseline PG-DPO algorithm, these costate estimates themselves (and their potential derivatives w.r.t. states) are not directly used in the parameter update rule (Step 6), which only requires the final gradient $\nabla_{(\theta,\phi)}J^{(i)}(\dots)$.
Nevertheless, the fact that BPTT provides access to these estimates is fundamental. Firstly, they are essential inputs for the Two-Stage PG-DPO variant discussed in the next subsection (§3.2). The baseline algorithm thus lays the groundwork by generating these estimates. Secondly, this connection explains why the algorithm is termed Pontryagin-Guided. Even though the baseline update only uses $\nabla J^{(i)}(\dots)$, the calculation of this gradient via BPTT structurally mirrors the backward propagation of sensitivities inherent in PMP’s adjoint equations, using the costate estimates $\lambda_{k}^{(i)},\bm{\eta}_{k}^{(i)}$ implicitly within the chain rule (cf. Huh, , 2024). Therefore, the optimization process, driven by these BPTT-computed gradients, is implicitly guided by the optimality conditions derived from Pontryagin’s principle, aiming to maximize the expected objective $\widetilde{J}$.
@@ 3.2 Two-Stage PG-DPO Algorithm
[…]
    $+\underbrace{\bm{\sigma}(t_{k},\mathbf{Y}_{k})^{\top}\!\bigl{[}(\partial_{\mathbf{Y}}\lambda_{k})\,\bm{\sigma}_{Y}(t_{k},\mathbf{Y}_{k})\,\rho^{\top}\,\Psi^{-1}\bigr{]}}_{\text{Hedging Demand}}\Biggr{\}}.$

### v2

@@ 3.1 Baseline PG-DPO for the Extended Objective
Algorithm 1 Baseline Pontryagin‑Guided Direct Policy Optimization (PG‑DPO)
1:Inputs:
2:Policy nets $\pi_{\theta},\,C_{\phi}$; learning‑rate schedule $\{\alpha_{j}\}_{j=0}^{J_{\max}-1}$; max iters $J_{\max}$, batch size $M$, time steps $N$; initial‑state distribution $\eta$; utility $U$, discount $\delta$, terminal weight $\kappa$; market params $(\bm{\mu},\bm{\sigma},r,\bm{\mu}_{Y},\bm{\sigma}_{Y},\bm{\Omega})$
3:Output: optimised params $(\theta_{\text{opt}},\phi_{\text{opt}})$
4:Initialise $\theta,\phi$
5:for $j=0$ to $J_{\max}-1$ do
6:     Draw mini‑batch $\{(t_{0}^{(i)},x_{0}^{(i)},\mathbf{y}_{0}^{(i)})\}_{i=1}^{M}\sim\eta$
7:     $\nabla_{\text{batch}}\leftarrow\mathbf{0}$
8:     for $i=1$ to $M$ do
9:         $(X_{0},\mathbf{Y}_{0})\leftarrow(x_{0}^{(i)},\mathbf{y}_{0}^{(i)})$, $\Delta t\leftarrow(T-t_{0}^{(i)})/N$
10:         for $k=0$ to $N-1$ do
11:              $t_{k}\leftarrow t_{0}^{(i)}+k\Delta t$
12:              $\bm{\pi}_{k}\leftarrow\pi_{\theta}(t_{k},X_{k},\mathbf{Y}_{k})$
13:              $C_{k}\leftarrow C_{\phi}(t_{k},X_{k},\mathbf{Y}_{k})$
14:              Sample $\Delta\mathbf{W}_{k}\sim\mathcal{N}(\mathbf{0},\bm{\Omega}\Delta t)$
15:              $\mathbf{Y}_{k+1}=\mathbf{Y}_{k}+\bm{\mu}_{Y}(t_{k},\mathbf{Y}_{k})\Delta t+\bm{\sigma}_{Y}(t_{k},\mathbf{Y}_{k})\,\Delta\mathbf{W}_{k}^{\text{Y}}$
16:              $X_{k+1}=X_{k}+\Bigl{[}X_{k}\bigl{(}r(t_{k},\mathbf{Y}_{k})+\bm{\pi}_{k}^{\!\top}(\bm{\mu}(t_{k},\mathbf{Y}_{k})-r(t_{k},\mathbf{Y}_{k})\bm{1})\bigr{)}-C_{k}\Bigr{]}\Delta t+X_{k}\bm{\pi}_{k}^{\!\top}\bm{\sigma}(t_{k},\mathbf{Y}_{k})\,\Delta\mathbf{W}_{k}^{\text{X}}$
17:         end for
18:         $J^{(i)}=\sum_{k=0}^{N-1}e^{-\delta t_{k}}U(C_{k})\Delta t+\kappa e^{-\delta T}U(X_{N})$
19:         $g^{(i)}\leftarrow\nabla_{(\theta,\phi)}J^{(i)}$
20:         $\nabla_{\text{batch}}\leftarrow\nabla_{\text{batch}}+g^{(i)}$
21:     end for
22:     $(\theta,\phi)\leftarrow(\theta,\phi)+\dfrac{\alpha_{j}}{M}\nabla_{\text{batch}}$
23:end for
24:return $(\theta,\phi)$
@@
It is important to understand the role of the intermediate values computed during BPTT in Step 4. As BPTT calculates the policy gradient $\nabla_{(\theta,\phi)}J^{(i)}(\dots)$ via the chain rule, it inherently computes the partial derivatives of the realized reward $J^{(i)}(\dots)$ with respect to the intermediate states along path $i$:
@@
These sensitivities, computed mechanically by AD, are directly linked to the discrete adjoint Backward Stochastic Differential Equation (BSDE) structure required by Pontryagin’s Maximum Principle (PMP). To illustrate this crucial connection for the wealth component (denoting $\lambda_{k}^{(i)}$ for $\lambda_{k}^{X,(i)}$ for simplicity), consider the discrete Euler-Maruyama step for wealth under policy $(\bm{\pi}_{k}^{(i)},C_{k}^{(i)})$:
$X_{k+1}^{(i)}=X_{k}^{(i)}+\Bigl{[}X_{k}^{(i)}\Bigl{(}r(t_{k},\mathbf{Y}_{k}^{(i)})+\bm{\pi}_{k}^{(i)\top}\bigl{(}\bm{\mu}(t_{k},\mathbf{Y}_{k}^{(i)})-r(t_{k},\mathbf{Y}_{k}^{(i)})\mathbf{1}\bigr{)}\Bigr{)}-C_{k}^{(i)}\Bigr{]}\Delta t+X_{k}^{(i)}\bm{\pi}_{k}^{(i)\top}\bm{\sigma}(t_{k},\mathbf{Y}_{k}^{(i)})\,\Delta\mathbf{W}_{k}^{(i)}.$
Applying the chain rule one step backward ($\lambda_{k}^{(i)}=\lambda_{k+1}^{(i)}\frac{\partial X_{k+1}^{(i)}}{\partial X_{k}^{(i)}}$) and using a first-order expansion yields the following backward recurrence relation along the path $(i)$:
$\lambda_{k}^{(i)}\approx\lambda_{k+1}^{(i)}+\lambda_{k+1}^{(i)}\bigl{[}r+\pi_{k}^{(i)}(\mu-r)\bigr{]}\Delta t+\lambda_{k+1}^{(i)}\pi_{k}^{(i)}\sigma\,\Delta W_{k}^{(i)}.$
This equation inherently possesses the structure of a discrete BSDE operating along the specific simulation path $(i)$. To formally connect this to the BSDE formulation in PMP, which involves processes adapted to the filtration $\{\mathcal{F}_{t_{k}}\}$ (representing available information at time $t_{k}$), one defines the adapted costate $\lambda_{k}=\mathbb{E}[\lambda_{k}^{(i)}|\mathcal{F}_{t_{k}}]$ and the predictable martingale component $Z_{k}$ (related to the conditional expectation of $\lambda_{k+1}^{(i)}\Delta W_{k}^{(i)}$). Taking the conditional expectation ($\mathbb{E}[\cdot|\mathcal{F}_{t_{k}}]$) of the pathwise recurrence relation above, and carefully defining $Z_{k}$ using a covariation-based approach, leads precisely to the standard discrete BSDE for the adapted process $\lambda_{k}$:
$\lambda_{k}=\lambda_{k+1}+\underbrace{\bigl{(}\lambda_{k+1}[r+\pi_{k}(\mu-r)]+\pi_{k}\,\sigma\,Z_{k}\bigr{)}}_{\approx\partial_{x}\mathcal{H}(t_{k},\dots)\text{ from PMP}}\Delta t-Z_{k}\,\Delta W_{k}.$ (13)
Crucially, the drift term derived through this process corresponds exactly to the derivative of the discrete-time Hamiltonian ($\partial_{x}\mathcal{H}$) from PMP with respect to the state variable $x$. A similar structure holds for the costate vector $\bm{\eta}_{k}$ associated with the state variables $\mathbf{Y}_{k}$.
Therefore, the values $\lambda_{k}^{(i)}$ and $\bm{\eta}_{k}^{(i)}$ computed by BPTT are indeed pathwise numerical counterparts related to the Pontryagin costates, satisfying the relevant discrete BSDE structure. Crucially, in the baseline PG-DPO algorithm, these costate estimates themselves (and their potential derivatives w.r.t. states) are not directly used in the parameter update rule (Step 6), which only requires the final gradient $\nabla_{(\theta,\phi)}J^{(i)}(\dots)$.
Nevertheless, the fact that BPTT provides access to these theoretically grounded estimates is fundamental. Firstly, they are essential inputs for the Two-Stage PG-DPO variant discussed in the next subsection (§3.2). The baseline algorithm thus lays the groundwork by generating these estimates. Secondly, this deep connection explains why the algorithm is termed Pontryagin-Guided. Even though the baseline update only uses $\nabla_{(\theta,\phi)}J^{(i)}(\dots)$, the calculation of this gradient via BPTT structurally mirrors the backward propagation of sensitivities inherent in PMP’s adjoint equations (BSDEs), using the costate estimates $\lambda_{k}^{(i)},\bm{\eta}_{k}^{(i)}$ implicitly within the chain rule (cf. Huh, , 2024). Consequently, the optimization process, driven by these BPTT-computed gradients, is implicitly guided by the optimality conditions derived from Pontryagin’s principle, aiming to maximize the expected objective $\widetilde{J}$.
@@ 3.2 Two-Stage PG-DPO Algorithm
Algorithm 2 Two‑Stage PG‑DPO: Warm‑up Training & Deployment
1:Stage 1: Warm‑up Training
2:Inputs of Algorithm 1; warm‑up iterations $K_{0}$
3:Stabilised parameters $(\theta^{*},\phi^{*})$
4:Run Algorithm 1 for $K_{0}$ iterations to obtain $(\theta^{*},\phi^{*})$
5:
6:Stage 2: Deployment at state $(t,X,\bm{Y})$
7:State $(t,X,\bm{Y})$; parameters $(\theta^{*},\phi^{*})$; Monte‑Carlo rollouts $M_{\mathrm{MC}}$, steps $N^{\prime}$
8:Near‑optimal controls $(\bm{\pi}^{\mathrm{PMP}},C^{\mathrm{PMP}})$
9:Initialise $\mathcal{L},\mathcal{L}_{x},\mathcal{L}_{Y}\leftarrow\varnothing$
10:for $j=1$ to $M_{\mathrm{MC}}$ do
11:     Simulate path with $(\pi_{\theta^{*}},C_{\phi^{*}})$ using $N^{\prime}$ steps
12:     Compute realized reward for path $j$:
13:     $\qquad\displaystyle J^{(j)}\leftarrow\sum_{n=0}^{N^{\prime}-1}e^{-\delta t_{n}}U\bigl{(}C_{t_{n}}^{(j)}\bigr{)}\,\Delta t^{\prime}+\kappa e^{-\delta T}U\bigl{(}X_{T}^{(j)}\bigr{)}$
14:     Evaluate $\lambda_{t}^{(j)}$, $\partial_{x}\lambda_{t}^{(j)}$, $\partial_{\bm{Y}}\lambda_{t}^{(j)}$ and append
15:end for
16:Compute $\hat{\lambda}_{t},\widehat{\partial_{x}\lambda}_{t},\widehat{\partial_{\bm{Y}}\lambda}_{t}$triangleright E.g., by averaging estimates from loop
17:$C^{\mathrm{PMP}}\leftarrow(U^{\prime})^{-1}\!\bigl{(}e^{\delta t}\hat{\lambda}_{t}\bigr{)}$triangleright Using PMP FOC Eq. (8)
18:Compute $\bm{\pi}^{\mathrm{PMP}}$ via PMP FOC Eq. (15) $\triangleright$ Using estimated costates/derivatives
19:return $(\bm{\pi}^{\mathrm{PMP}},C^{\mathrm{PMP}})$
@@
    $\bm{\pi}_{k}^{\mathrm{PMP}}$ $=-\frac{1}{X_{k}\,\partial_{x}\lambda_{k}}\,\Sigma^{-1}\!\bigl{(}t_{k},\mathbf{Y}_{k}\bigr{)}\Bigl{\{}\lambda_{k}\bigl{[}\bm{\mu}(t_{k},\mathbf{Y}_{k})-r(t_{k},\mathbf{Y}_{k})\mathbf{1}\bigr{]}$ (14)
    $\qquad+\bm{\sigma}(t_{k},\mathbf{Y}_{k})^{\top}(\partial_{\mathbf{Y}}\lambda_{k})\,\bm{\sigma}_{Y}(t_{k},\mathbf{Y}_{k})\,\rho^{\top}\Psi^{-1}\Bigr{\}}.$ (15)